(D^2+2DD'+D')=e^x-y+xy
1
Expert's answer
2022-01-24T14:18:24-0500

auxillary equation:

"m^2+2m+1=0"

"m_{1,2}=-1"

"C.F.=f_1(y+mx)+xf_2(y+mx)=f_1(y-x)+xf_2(y-x)"


for particular integral:

for "e^{x-y}" :

"\frac{1}{F(D,D')}e^{ax+by}=\frac{1}{F(a,b)}e^{ax+by}"


"\frac{1}{D^2+2DD'+D'}e^{x-y}=\frac{1}{1-2-1}e^{x-y}=-e^{x-y}/2"


for "xy" :

"\frac{1}{F(D,D')}xy=\frac{1}{D^2}(1+\frac{2D'}{D}+\frac{D'}{D^2})^{-1}xy=\frac{1}{D^2}(1-\frac{2D'}{D}+...)xy="


"=\frac{1}{D^2}(xy-2x/D)=x^3/6-x^4/12"


"z=f_1(y-x)+xf_2(y-x)-e^{x-y}\/2+x^3\/6-x^4\/12"